Find the amount of an ordinary annuity in which you make a payment of $1500 at the end of 3 months at 8% per annum, compounded quarterly.

I am having trouble with this problem because I feel as though there is missing information (specifically, the length of time over which these payments are made or how long the annuity is to continue) but the textbook provides more than one question of this type, so I assume that this was intentional. I would appreciate an explanation as soon as possible.

Thank You!!!
Answer #1

PV = Periodic Payment / Rate of Interest of periodic time period.

Rate of Interest of periodic time period. = 0.08 / 4 = 0.02

PV = 1500 / 0.02

PV = 75000

here, the case is of perpetuity. 1500$ will be received forever.

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.